A six-month-old child is on a maintenance dose of Digoxin ordered at 0.04 mg. orally twice daily. The child's weight is 14 1/2 lb. If the order is safe, how many mLs would be given with one dose?
Answer #1

Child's weight 14 1/2lb= 6.57 kg

Dose :0.04 mg twice daily = 0.04*2= 0.08mg/day

Normal maintenance dose is 0.01-0.15 mg/kg

6.57*0.01mg/ kg=0.0657 minimum dosage

6.57 kg*0.15mg/kg=0.985 Maximum dosage

The drug is within range

0.08 mg/day is within range

0.08 ml would be given with one dose